# Bounds on the number of squares in recurrence sequences

### v1

@@ 4.6. Small $d$
(1) We start by considering $t^{2}-du^{2}=4$.
(1-i) If $u=4$ and $t^{2}-du^{2}=4$, then we can argue as for $u=8$ in the proof of Lemma 5.3 and use the Maple program mentioned there modulo $9$ to show that $9|\gcd\left(a^{2},d\right)$. Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 144$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq(76/144)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}>4\left|N_{\alpha}\right|$ here. So we can exclude $u=4$ and $t^{2}-du^{2}=4$ from consideration.
(1-ii) Suppose that $u=3$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=4$. We will again argue as for $u=8$ in the proof of Lemma 5.3 and use the Maple program mentioned there modulo $32$ to show that $64|\gcd\left(a^{2},d\right)$.
Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 256$. Hence the theorem holds if $y_{\pm 1}>(76/256)\left|N_{\alpha}\right|$, by Proposition 4.1. But by Lemma 3.5(c), we know that $y_{\pm 1}>(9/4)\left|N_{\alpha}\right|$ here.
(1-iii) Suppose that $u=2$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=4$. Using the argument for $u=8$ in the proof of Lemma 5.3 and the Maple program mentioned there modulo $16$, we obtain $16|\gcd\left(a^{2},d\right)$.
Doing the same modulo $9$, we obtain $9|\gcd\left(a^{2},d\right)$. Combining this with $16|\gcd\left(a^{2},d\right)$ and applying Lemma 3.11, we have $|g|\mathcal{N}_{d,4}\geq 24$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq\left(76/24^{2}\right)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}>\left|N_{\alpha}\right|u^{2}/4=\left|N_{\alpha}\right|$ here. So we can exclude $u=2$ and $t^{2}-du^{2}=4$ from consideration.
(1-iv) Suppose that $u=1$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=4$. Here too we use the argument and Maple program from the proof of Lemma 5.3, first modulo $9$ and then modulo $16$. From the latter, we obtain $64|\gcd\left(a^{2},d\right)$. So by Lemma 3.11, $|g|\mathcal{N}_{d,4}\geq 48$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq\left(76/48^{2}\right)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}>\left|N_{\alpha}\right|/4$ here. So we can exclude $u=1$ and $t^{2}-du^{2}=4$ from consideration.
(2) We now suppose that $t^{2}-du^{2}=-4$.
@@
(2-i) Consider $u=1$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=-4$.
@@
If $8|b_{1}$, then $a\equiv 2\pmod{4}$ and $|g|\mathcal{N}_{d,4}\geq 16$. So $\frac{76\left|N_{\alpha}\right|^{3/2}}{\left(\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\right)}\leq\left(76/16^{2}\right)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$. However $76/16^{2}$ is bigger than $1/4$, so we must work a bit harder to eliminate $8|b_{1}$.
@@
Now we show that $\gcd\left(a^{2},d\right)=1$ or $4$.
@@
We consider first $\gcd\left(a^{2},d\right)>4$ odd. So by Lemma 3.11 and the argument in the previous paragraph, $|g|\mathcal{N}_{d,4}=4\sqrt{\gcd\left(a^{2},d\right)}$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{16\sqrt{d}\gcd\left(a^{2},d\right)}$. If $\gcd\left(a^{2},d\right)\geq 25$, then the right-hand side is less than $40.19\left|N_{\alpha}\right|<\left|N_{\alpha}\right|/$ 4, the right-hand side being the lower bound for $y_{\pm 1}$ from Lemma 3.5(c).
@@
Now consider $\gcd\left(a^{2},d\right)>4$ even. Then $v_{2}\left(\gcd\left(a^{2},d\right)\right)=2$ (so $\gcd\left(a^{2},d\right)\equiv 4\pmod{16}$). So by Lemma 3.11 and the argument above, $|g|\mathcal{N}_{d,4}=2\sqrt{\gcd\left(a^{2},d\right)}$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{4\sqrt{d}\gcd\left(a^{2},d\right)}$.
@@
(2-ii) Lastly, we consider $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$. Arguing modulo $64$, we have $d\equiv a^{2}\equiv 1\pmod{16}$ and $8|t$. So $16|N_{\alpha}$ and, in the notation of Lemma 3.11, $16|b^{\prime}$. Thus $|g|\mathcal{N}_{d^{\prime},4}=8\sqrt{\gcd\left(a^{2},d\right)}$, where $\gcd\left(a^{2},d\right)$ is odd.
@@
Theorem 1.4 now follows from the bounds on $u$ from Proposition 4.1, along with Lemmas 5.1, lem:nrmA-odd and lem:nrmA-even.
@@ 6. Examples
1 $3^{2}$ [0, -1, 3, -5] [3, 5, 31, 167]
1019 $27^{2}$ [0, 1, -3, -7] [27, 65, 29, 983]
167 $13^{2}$ [0, 1, -3, 4] [13,29,71, 407]
157 $29^{2}$ [0, -1, 3, -4] [29, 47, 307, 649]
1 $41^{2}$ [0, -1, -9, 11] [41, 71, 80753, 470861]
1633 $65^{2}$ [0, -1, -4, 7] [65, 97, 1331, 24791]
48479 $211^{2}$ [0, -3, 4, -7] [211, 1007, 6743, 34205]
45649 $677^{2}$ [0, -1, -4, 4] [677, 1133, 15679, 16825]
1940147 $1217^{2}$ [0, -3, 4, -11] [1217, 3289, 40573, 3794239]
600589 $2213^{2}$ [0, -1, -4, 4] [2213, 3673, 50801, 55415]
20509501 $8689^{2}$ [0, -1, 3, -4] [8689, 13619, 94393, 187603]
255488029 $13457^{2}$ [0, -1, 3, -4] [13457, 5683, 189241, 15821]
409660129 $17023^{2}$ [0, -1, -4, -8] [17023, 7073, 7949, 269495]
3032771269 $46313^{2}$ [0, -1, -4, -8] [46313, 19213, 15269, 516625]
We also found such examples with $(d,t,u,a,b)=(3,4,2,672,8281)$, $(6,10,4,78,49)$, $(6,10,4,34986,22201)$, $(6,10,4,3663828,5094049)$, $(30,22,4,826320,1234321)$ and $(37,12,2,138,25)$.

### v2

@@ 4.6. Small $d$
(1) We start by showing that $t^{2}-du^{2}=4$ is not possible.
(1-i) We show that $u=4$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
Here we can argue as for $u=8$ in the proof of Lemma 5.5 and use the Maple program mentioned there modulo $9$ to show that $9|\gcd\left(a^{2},d\right)$. Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 144$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq(76/144)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}\geq 4\left|N_{\alpha}\right|$ here. So we can exclude $u=4$ and $t^{2}-du^{2}=4$ from consideration.
(1-ii) We show that $u=3$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
We will again argue as for $u=8$ in the proof of Lemma 5.5 and use the Maple program mentioned there modulo $32$ to show that $64|\gcd\left(a^{2},d\right)$.
Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 256$. Hence the theorem holds if $y_{\pm 1}>(76/256)\left|N_{\alpha}\right|$, by Proposition 4.1. But by Lemma 3.5(c), we know that $y_{\pm 1}\geq(9/4)\left|N_{\alpha}\right|$ here.
(1-iii) We show that $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
Using the argument for $u=8$ in the proof of Lemma 5.5 and the Maple program mentioned there modulo $16$, we obtain $16|\gcd\left(a^{2},d\right)$.
Doing the same modulo $9$, we obtain $9|\gcd\left(a^{2},d\right)$. Combining this with $16|\gcd\left(a^{2},d\right)$ and applying Lemma 3.12, we have $|g|\mathcal{N}_{d^{\prime},4}\geq 24$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq\left(76/24^{2}\right)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}\geq\left|N_{\alpha}\right|u^{2}/4=\left|N_{\alpha}\right|$ here. So we can exclude $u=2$ and $t^{2}-du^{2}=4$ from consideration.
(1-iv) We show that $u=1$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=4$ is not possible.
Here too we use the argument and Maple program from the proof of Lemma 5.5, first modulo $9$ and then modulo $16$. From the latter, we obtain $64|\gcd\left(a^{2},d\right)$. So by Lemma 3.12, $|g|\mathcal{N}_{d^{\prime},4}\geq 48$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\leq\left(76/48^{2}\right)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}\geq\left|N_{\alpha}\right|/4$ here. So we can exclude $u=1$ and $t^{2}-du^{2}=4$ from consideration.
(2) We now consider $t^{2}-du^{2}=-4$.
@@
(2-i) We show that if $u=1$, $-N_{\alpha}$ is an even square and $t^{2}-du^{2}=-4$, then $N_{\alpha}\equiv 12\pmod{16}$.
@@
If $8|b_{1}$, then $a\equiv 2\pmod{4}$ and $|g|\mathcal{N}_{d^{\prime},4}\geq 16$. So $\frac{76\left|N_{\alpha}\right|^{3/2}}{\left(\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\right)}\leq\left(76/16^{2}\right)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$. However $76/16^{2}$ is bigger than $1/4$, so we must work a bit harder to eliminate $8|b_{1}$.
@@
(2-ii) We show that if $u=1$, $-N_{\alpha}$ is an even square, $t^{2}-du^{2}=-4$, and $N_{\alpha}\equiv 12\pmod{16}$, then $\gcd\left(a^{2},d\right)=1$ or $4$.
@@
We consider first $\gcd\left(a^{2},d\right)>4$ odd. So by Lemma 3.12 and the argument in the previous paragraph, $|g|\mathcal{N}_{d^{\prime},4}=4\sqrt{\gcd\left(a^{2},d\right)}$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{16\sqrt{d}\gcd\left(a^{2},d\right)}$. If $\gcd\left(a^{2},d\right)\geq 25$, then the right-hand side is less than $40.19\left|N_{\alpha}\right|<\left|N_{\alpha}\right|/$ 4, the right-hand side being the lower bound for $y_{\pm 1}$ from Lemma 3.5(c).
@@
Now consider $\gcd\left(a^{2},d\right)>4$ even. Then $v_{2}\left(\gcd\left(a^{2},d\right)\right)=2$ (so $\gcd\left(a^{2},d\right)\equiv 4\pmod{16}$). So by Lemma 3.12 and the argument above, $|g|\mathcal{N}_{d^{\prime},4}=2\sqrt{\gcd\left(a^{2},d\right)}$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}=\frac{76\left|N_{\alpha}\right|^{3/2}}{4\sqrt{d}\gcd\left(a^{2},d\right)}$.
@@
(2-iii) We show that $u=2$, $-N_{\alpha}$ an even square and $t^{2}-du^{2}=-4$ is not possible.
Arguing modulo $64$, we have $d\equiv a^{2}\equiv 1\pmod{16}$ and $8|t$. So $16|N_{\alpha}$ and, in the notation of Lemma 3.12, $16|b^{\prime}$. Thus $|g|\mathcal{N}_{d^{\prime},4}=8\sqrt{\gcd\left(a^{2},d\right)}$, where $\gcd\left(a^{2},d\right)$ is odd.
@@
Theorem 1.4(c) now follows from the bounds on $u$ that we obtained from Proposition 4.1 at the start of Section 5, along with Lemmas 5.1, 5.2, 5.5 and 5.6. Furthermore, from Lemma 5.3, if $y_{-1}$ and $y_{1}$ are both squares, then there are no further squares. So we may assume that precisely one of $y_{\pm 1}$ is a square. From Lemma 3.5(c), we obtain $y_{k}>19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$ for $|k|\geq 2$, provided that $d>190$ when $u=1$ and that $d>11$ when $u=2$. This leaves $26$ values of $d$. For each of these, we compute $y_{-2}$ directly for each possibility of $N_{\alpha}$ and compare the value to the bound in Proposition 4.1. Where the bound was exceeded, Proposition 4.1 tells us that there are no further squares, completing the proof for them. There were only four cases where $y_{-2}$ did not exceed that bound: $(a,b,d,t,u)=(1,1,2,2,2)$, $(1,1,5,1,1)$, $(2,1,8,2,1)$, $(3,1,13,3,1)$. These were treated in Subsection 4.6.
@@ 6. Examples
1 $3$ [0, -1, 3, -5] [3, 5, 31, 167]
1019 $27$ [0, 1, -3, -7] [27, 65, 29, 983]
167 $13$ [0, 1, -3, 4] [13,29,71, 407]
157 $29$ [0, -1, 3, -4] [29, 47, 307, 649]
1 $41$ [0, -1, -9, 11] [41, 71, 80753, 470861]
1633 $65$ [0, -1, -4, 7] [65, 97, 1331, 24791]
48479 $211$ [0, -3, 4, -7] [211, 1007, 6743, 34205]
45649 $677$ [0, -1, -4, 4] [677, 1133, 15679, 16825]
1940147 $1217$ [0, -3, 4, -11] [1217, 3289, 40573, 3794239]
600589 $2213$ [0, -1, -4, 4] [2213, 3673, 50801, 55415]
20509501 $8689$ [0, -1, 3, -4] [8689, 13619, 94393, 187603]
255488029 $13457$ [0, -1, 3, -4] [13457, 5683, 189241, 15821]
409660129 $17023$ [0, -1, -4, -8] [17023, 7073, 7949, 269495]
3032771269 $46313$ [0, -1, -4, -8] [46313, 19213, 15269, 516625]
Table 1. Examples for $d=2$ with $t=u=2$
In addition to the examples for $d=2$ in Table 1, we also found examples with four squares for $(d,t,u,a,b)=(3,4,2,672,91)$, $(6,10,4,78,7)$, $(6,10,4,34986,149)$,
$(6,10,4,3663828,2257)$, $(30,22,4,826320,1111)$ and $(37,12,2,138,5)$.
@@ 1 $41$ [0, -1, -9, 11] [41, 71, 80753, 470861]
Here are two examples showing that there are at least three distinct squares when $\left|N_{\alpha}\right|$ is a square:
$(d,t,u,a,b)=(5,1,1,43,3)$: $N_{\alpha}=38^{2}$, $y_{-3}=5^{2}$, $y_{0}=3^{2}$, $y_{11}=53^{2}$;
$(d,t,u,a,b)=(10,6,2,1,1)$: $N_{\alpha}=-3^{2}$, $y_{0}=1^{2}$, $y_{1}=2^{2}$, $y_{2}=5^{2}$.
Here are two examples showing that the same is true when $\left|N_{\alpha}\right|$ is a prime power:
$(d,t,u,a,b)=(5,1,1,153,4)$: $N_{\alpha}=22129$, $y_{-3}=11^{2}$, $y_{0}=4^{2}$, $y_{30}=8862^{2}$;
$(d,t,u,a,b)=(51,100,14,2,1)$: $N_{\alpha}=-47$, $y_{-1}=6^{2}$, $y_{0}=1^{2}$, $y_{1}=8^{2}$.
Here is an example showing that if $\left|N_{\alpha}\right|$ is not a prime power, but $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)$ is a prime, then there can be distinct four squares (unlike in Conjecture 1.1):
$(d,t,u,a,b)=(5,1,1,7,1)$: $N_{\alpha}=2^{2}\cdot 11$, $y_{-9}=9^{2}$, $y_{0}=1$, $y_{1}=2^{2}$, $y_{3}=3^{2}$.
And finally, an example showing that if $\operatorname{sf}\left(\left|N_{\alpha}\right|\right)$ is twice an odd prime, then there can be distinct four squares (again unlike in Conjecture 1.1):
$(d,t,u,a,b)=(6,10,4,2,3)$: $N_{\alpha}=-2\cdot 241$, $y_{-3}=63^{2}$, $y_{0}=3^{2}$, $y_{1}=7^{2}$, $y_{3}=69^{2}$.
6.4. Examples for Theorem 1.4(a)
From (3.14) with $b=u=1$, we find that $(2t+a)^{2}-8y_{1}=a^{2}-8$. Since $-N_{\alpha}=d-a^{2}=t^{2}+4-a^{2}$ is a square, we put $a=2$, so if $y_{1}$ is a square, then $22t+$ 2 is an element of the sequence $\left(t_{n}\right)_{n\geq 0}$ with $t_{0}=6$, $t_{1}=238$ and $t_{n+1}=34u_{n}-u_{n-1}$. The corresponding values of $\sqrt{y_{1}}$ are the elements of the sequence $\left(u_{n}\right)_{n\geq 0}$ with $u_{0}=5$, $u_{1}=169$ and $u_{n+1}=34u_{n}-u_{n-1}$. Then $y_{1}$ is approximately $t^{2}/2\approx\left|N_{\alpha}\right|/2\approx\left|N_{\alpha}\right|^{3/2}/\left(2\sqrt{d}\right)$.
In fact, such $y_{1}$’s are the ones that are squares satisfying the conditions in Theorem 1.4(a) with the smallest ratio compared to $\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$, the quantity in Proposition 4.1.
@@
The sequences given in Lemma 3.6 show that for $k=-1$, the lower bound in Lemma 3.5(c) can be attained too. At least, when the lower bound is $1$.